OpenStudy (anonymous):

Prove the identity. 2cos^2 x/1-cos^2 x = 2/tan^2 x

hartnn (hartnn):

do you recall a trigonometric identity relating sin^2 x and cos^2 x ?

OpenStudy (anonymous):

yes i do, i tried to sub in what fits to the identities but I'm still messing up

hartnn (hartnn):

ok, so we have \(\Large \sin^2 x+\cos^2 x=1\) try to isolate sin^2 x ?

hartnn (hartnn):

subtract cos^2 x from both sides, what u get ?

OpenStudy (anonymous):

so like this? sin^2x - cos^2 x = 1-cos^2 x

hartnn (hartnn):

\(\Large \sin^2 x+\cos^2 x=1\) \(\Large \sin^2 x+\cos^2 x-\cos^2 x=1-\cos^2 x \) \(\Large \sin^2 x = 1-\cos^2 x\) got this ?

OpenStudy (anonymous):

oh yeah i forgot the first cos^2 but yup i get it :)

hartnn (hartnn):

so your denominator is just sin^2 x, right ?

OpenStudy (anonymous):

yup :) and now from there i have to make it equal to the 2/tan^2 x?

hartnn (hartnn):

correct

OpenStudy (anonymous):

well that was easy, thank you! :)

hartnn (hartnn):

welcome :)